# APCTP-Pre2023-010

### v2

@@ II.1 Orthonormal frame and initial conditions
$\ddot{q}^{a}+\left(3H\delta^{a}_{~{}b}-2z^{a}_{~{}b}\right)\dot{q}^{b}+\left({k^{2}\over a^{2}}\delta^{a}_{~{}b}+M_{IJ}E^{I}_{~{}a}E^{J}_{~{}b}+z^{a}_{~{}c}z^{c}_{~{}b}-3Hz^{a}_{~{}b}-\dot{z}^{a}_{~{}b}\right)q^{b}=0\,.$ (16)
By an appropriate field redefinition, and using the Frenet-Serret equations, we can simplify the Lagrangian (15). By redefining the tangential component as $q_{\rm t}\equiv\sqrt{2\epsilon}\mathcal{R}$ and performing integrations by parts, the Lagrangian simplifies to
$L$ $=a^{3}\left\{{1\over 2}\dot{q}^{a}\dot{q}_{a}-{1\over 2a^{2}}\nabla q^{a}\nabla q_{a}-{\epsilon\over a^{2}}\nabla\mathcal{R}\nabla\mathcal{R}-z_{ab}\dot{q}^{a}q^{b}+{1\over 2}\omega^{2}q_{\rm n}^{2}+\epsilon\dot{\mathcal{R}}^{2}-2\omega\sqrt{2\epsilon}\dot{\mathcal{R}}q_{\rm n}-{1\over 2}\left(M_{ab}+z_{ac}z^{c}_{~{}b}\right)q^{a}q^{b}\right.$ (17)
$\quad\left.+\left(M_{\mathcal{R}{\rm b}}+\omega\tau_{\rm b}\right)\sqrt{2\epsilon}\mathcal{R}q_{\rm b}-\left[M_{\mathcal{R}\mathcal{R}}-\omega^{2}-{1\over 4}\eta^{2}H^{2}-{3\over 2}H\dot{\epsilon}-{1\over\epsilon}{\,\mathrm{d}\over\,\mathrm{d}t}\left({\epsilon\eta H\over 2}\right)\right]\epsilon\mathcal{R}^{2}+\left(M_{\mathcal{R}{\rm n}}+3H\omega+\dot{\omega}+{\dot{\epsilon}H\over\epsilon}\right)\sqrt{2\epsilon}\mathcal{R}q_{\rm n}\right\}\,,$
where now $a,b\neq\mathcal{R}$ and $\eta\equiv\dot{\epsilon}/(H\epsilon)$. Using the Frenet-Serret equations we find that the second line which involves $M_{\mathcal{R}\mathcal{R}}$, $M_{\mathcal{R}\rm{n}}$ and $M_{\mathcal{R}\rm{b}}$ vanishes identically. Thus, the quadratic Lagrangian for multi-field perturbations in terms of the curvature perturbation $\mathcal{R}$ and the isocurvature perturbations $q_{i}$ is
$L=a^{3}\left[{1\over 2}\dot{q}^{i}\dot{q}_{i}-z_{ij}\dot{q}^{i}q^{j}-{1\over 2a^{2}}\nabla q^{i}\nabla q_{i}-{1\over 2}\left(M_{ij}+z_{im}z^{m}{}_{j}\right)q^{i}q^{j}+{1\over 2}\omega^{2}q_{\rm n}^{2}-2\omega\sqrt{2\epsilon}\dot{\mathcal{R}}q_{\rm n}-{\epsilon\over a^{2}}\nabla\mathcal{R}\nabla\mathcal{R}+\epsilon\dot{\mathcal{R}}^{2}\right]\,.$ (18)
Similar expressions for the second-order Lagrangian have been derived in Achucarro:2018ngj ; Pinol:2020kvw .
To properly assign initial conditions for the perturbations (e.g. Bunch-Davies) we need to transform the action into its canonical form. This can be done by using a transformation to another local orthonormal basis which can eliminate the kinetic couplings between the fields. We will work with the conformal time $\,\mathrm{d}\tau\equiv a^{-1}\,\mathrm{d}t$ and redefine the fields as $u^{i}\equiv aq^{i}$ to simplify the action (15) to
$S_{2}={1\over 2}\int\,\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\delta_{ij}{\,\mathrm{d}u^{i}\over\,\mathrm{d}\tau}{\,\mathrm{d}u^{j}\over\,\mathrm{d}\tau}-2aHZ_{ij}{\,\mathrm{d}u^{i}\over\,\mathrm{d}\tau}u^{j}-\delta_{ij}\nabla u^{i}\nabla u^{j}-(aH)^{2}m_{ij}u^{i}u^{j}\right]\,,$ (19)
@@
$\begin{split}Z_{ij}&\equiv\frac{z_{ij}}{H}\,,\\ \mathcal{M}_{ij}&\equiv\frac{M_{ij}}{H^{2}}\,,\\ m_{ij}&\equiv\mathcal{M}_{IJ}E_{i}^{~{}I}E_{j}^{~{}J}+Z_{il}Z^{l}_{~{}j}-\left(2-\epsilon\right)\delta_{ij}\,.\end{split}$ (20)
The expression (19) is not yet suitable for quantization due to the second term that mixes the fields and their derivatives. We can eliminate these terms by introducing explicit time dependence in the definition of the orthonormal basis vectors Achucarro:2010da ; Lalak:2007vi ; Cremonini:2010ua . A time-dependent change of basis
$u^{i}=R^{i}_{~{}\alpha}(\tau)v^{\alpha}\,,$ (21)
@@
${\,\mathrm{d}R^{i}_{~{}\alpha}\over\,\mathrm{d}\tau}-aHZ^{i}_{~{}j}R^{j}_{~{}\alpha}=0\,,$ (22)
i.e. if the matrix is orthogonal. Regarding the initial condition of (22), we are completely free to choose any convenient, physically relevant time. In accordance to the previous literature we consider that the two basis vectors are aligned at horizon crossing $k\tau_{*}=-1$ and hence the initial condition becomes $R^{i}_{~{}\alpha}(\tau_{*})=\delta^{i}{}_{\alpha}$. In terms of the new fields we find a canonical second-order action as
$S_{2}={1\over 2}\int\,\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\delta_{\alpha\beta}{\,\mathrm{d}v^{\alpha}\over\,\mathrm{d}\tau}{\,\mathrm{d}v^{\beta}\over\,\mathrm{d}\tau}-\delta_{\alpha\beta}\nabla v^{\alpha}\nabla v^{\beta}+(aH)^{2}(2-\epsilon)\delta_{\alpha\beta}v^{\alpha}v^{\beta}-(aH)^{2}\mathcal{M}_{ij}R^{i}_{~{}\alpha}R^{j}_{~{}\beta}v^{\alpha}v^{\beta}\right]\,,$ (23)
@@
$v^{\alpha}|_{\rm BD}={1\over\sqrt{2k}}e^{-ik\tau}\left(1-{i\over k\tau}\right)\,,\qquad\alpha=1,\cdots,\mathcal{N}\,,$ (25)
@@
$u^{i}|_{-k\tau\gg 1}=R^{i}_{~{}\alpha}(\tau,\tau_{*})v^{\alpha}|_{\rm BD}\,.$ (26)
II.2 Scale dependence of the power spectrum
On sufficiently late times we are interested in the scale dependence of the normalized power spectrum defined by
$\left\langle{\mathcal{R}}(\bm{k}){\mathcal{R}}(\bm{k}^{\prime})\right\rangle\equiv(2\pi)^{3}\delta^{(3)}\left(\bm{k}-\bm{k}^{\prime}\right){2\pi^{2}\over k^{3}}\mathcal{P}_{\mathcal{R}}\,.$ (27)
Due to the complexity of the coupled equations, in general one has to solve numerically the coupled equations of perturbations to find the superhorizon value of $\mathcal{R}$. However, we will show that when every background quantity, i.e. $\epsilon$, $\mathcal{M}_{\rm nn}\cdots$, is constant the scale dependence of the power spectrum can be read off the equations of perturbations without explicitly solving them. This can be done in the following way: First, in the differential equations for the independent variables of (24) we trade $\tau$ with Gong:2001he
$x\equiv-k\tau\,,$ (28)
@@ II.2 Scale dependence of the power spectrum
$\tilde{v}\tilde{v}_{,x}^{*}-\tilde{v}^{*}\tilde{v}_{,x}=-i\,.$ (30)
@@
$\mathcal{R}_{\rm as}={H\over\sqrt{2\epsilon k^{3}}}xR_{~{}\alpha}^{1}\tilde{v}^{\alpha}_{\rm as}\,,$ (31)
@@
$(1-\epsilon)N=-\log x+\log\left(\frac{k}{a_{0}H_{0}}\right)\,,$ (33)
@@
$H=H_{0}x^{\epsilon/(1-\epsilon)}\left({k\over a_{0}H_{0}}\right)^{-\epsilon/(1-\epsilon)}\,,$ (34)
and the normalized power spectrum scales as $\mathcal{P}_{\mathcal{R}}\sim k^{-2\epsilon/(1-\epsilon)}$, irrespective of the number of fields. This remarkably has exactly the same form as the single-field power spectrum for scaling solutions (see e.g. Stewart:1993bc for the calculation in the single-field case).
@@
$\mathcal{P}_{\mathcal{R}}={H^{2}\over 4\pi^{2}}\frac{x^{2}|\tilde{u}_{\rm t}|^{2}}{\epsilon}\,.$ (35)
@@
$|\tilde{u}_{\rm t}|^{2}\approx\tilde{u}_{i}^{*}\tilde{u}_{i}=\tilde{v}_{\alpha}^{*}\tilde{v}_{\alpha}\,,$ (36)
and the final result does not depend on the rotation matrix $R$. If $\epsilon$ and $\eta$ vary slowly, but all other quantities are constant then the scale dependence of the power spectrum becomes identical to the single-field case to lowest order. [^3] This has been noticed before in cases where the orthogonal perturbation is heavy and can be integrated out, leading to an effective description of the curvature perturbation with a modified speed of sound Achucarro:2010da . However, our conclusion is not limited to heavy orthogonal fields. Otherwise, to find the asymptotic value of the curvature perturbation one has to solve the differential equations explicitly, which is the topic of section III.2.
II.3 Hamiltonian description
@@
The canonical momenta for the Lagrangian (18) are
$\pi_{\mathcal{R}}$ $={\partial L\over\partial\dot{\mathcal{R}}}=2a^{3}\left(\epsilon\dot{\mathcal{R}}-\omega\sqrt{2\epsilon}q_{\rm n}\right)\,,$ (37)
$\pi_{i}$ $={\partial L\over\partial\dot{q}^{i}}=a^{3}\left(\dot{q}_{i}-z_{ij}q^{j}\right)\,,$ (38)
@@
$\mathcal{H}={1\over 2}{\pi^{i}\pi_{i}\over a^{3}}+z_{ij}\pi^{i}q^{j}+a{1\over 2}\nabla q^{i}\nabla q_{i}+{1\over 2}a^{3}M_{ij}q^{i}q^{j}+{3\over 2}a^{3}\omega^{2}q_{\rm n}^{2}+2{\omega\over\sqrt{2\epsilon}}\pi_{\mathcal{R}}q_{\rm n}+{1\over 4}{\pi_{\mathcal{R}}^{2}\over\epsilon a^{3}}+{\epsilon\over a^{2}}\nabla\mathcal{R}\nabla\mathcal{R}\,.$ (39)
@@
$\pi_{\mathcal{R}}$ $=a^{3}\sqrt{2\epsilon}\left(\,\text{D}_{t}q_{\rm t}-\frac{1}{2}H\eta q_{\rm t}-\omega q_{\rm n}\right)\,,$ (40)
$\pi_{i}$ $=a^{3}\left(\,\text{D}_{t}q_{i}-\omega q_{\rm n}\delta_{i{\rm n}}\right)\,.$ (41)
@@
$\dot{q}^{a}={\partial\mathcal{H}\over\partial\pi_{a}}\quad\text{and}\quad\dot{\pi}_{a}=-{\partial\mathcal{H}\over\partial q^{a}}$ (42)
@@
$\dot{\mathcal{R}}$ $={1\over 2}{\pi_{\mathcal{R}}\over\epsilon a^{3}}+\sqrt{\frac{2}{\epsilon}}\omega q_{\rm n}\,,$ (43)
$\dot{\pi}_{\mathcal{R}}$ $=2\epsilon a\nabla^{2}\mathcal{R}\,,$ (44)
$\dot{q}_{i}$ $={\pi_{i}\over a^{3}}+z_{ij}q^{j}\,,$ (45)
$\dot{\pi}_{i}$ $=-2{\omega\over\sqrt{2\epsilon}}\pi_{\mathcal{R}}\delta_{i{\rm n}}+z_{ij}\pi^{j}+a^{3}\left({1\over a^{2}}\nabla^{2}q_{i}-M_{ij}q^{j}-3\omega^{2}\delta_{i{\rm n}}q_{\rm n}\right)\,.$ (46)
@@
$P_{\cal R}\equiv\frac{\pi_{\cal R}}{2\epsilon a^{3}H}\quad\text{and}\quad P_{i}\equiv\frac{\pi_{i}}{a^{3}H}\,,$ (47)
@@
$\mathcal{R}^{\prime}$ $=P_{\mathcal{R}}+\sqrt{\frac{2}{\epsilon}}\Omega q_{\rm n}\,,$ (48)
$P_{\mathcal{R}}^{\prime}$ $=-(3-\epsilon+\eta)P_{\mathcal{R}}-\kappa^{2}\mathcal{R}\,,$ (49)
$q_{i}^{\prime}$ $=P_{i}+Z_{ij}q^{j}\,,$ (50)
$P_{i}^{\prime}$ $=-(3-\epsilon)P_{i}-2\sqrt{2\epsilon}\Omega P_{\mathcal{R}}\delta_{i{\rm n}}+Z_{ij}P^{j}-\kappa^{2}q_{i}-\mathcal{M}_{ij}q^{j}-3\Omega^{2}\delta_{i{\rm n}}q_{\rm n}\,,$ (51)
@@
$\Omega\equiv{\omega\over H}\quad\text{and}\quad\kappa\equiv{k\over aH}\,.$ (52)
The formulation in terms of the fields and their “momenta” [^4] Even though we will keep referring to the variables $P_{i}$ as momenta, these are not the canonical momenta of the problem. nicely illustrates the superhorizon sourcing of the curvature perturbation due to a non-vanishing normal perturbation Gordon:2000hv : Excluding runaway solutions $\mathcal{R}\propto a^{2}$ which would render the perturbation theory invalid, (49) implies that for $\kappa\rightarrow 0$ the momentum $P_{\mathcal{R}}$ is exponentially decaying. Therefore, $\mathcal{R}$ is sourced by the normal perturbation outside the horizon, whereas $q_{\rm n}$ is not. Moreover, for a sudden change of the bending parameters the previous set of equations has no singularities and no special junction conditions are needed at the discontinuities (as was done in e.g. Palma:2020ejf ; Fumagalli:2020nvq ; Boutivas:2022qtl ; Aragam:2023adu ) apart from continuity of the fields and their momenta.
III Analytic solutions for constant turns
To gain further physical insights, in this section we consider a simple two-field system that undergoes a sharp or broad turn. We can find a set of analytic solutions of perturbations, and in particular can identify how the positive eigenvalues enhance the perturbations. Note that while we use Hamilton’s equations to study the behaviour of the two perturbations, we have checked that our results for a single turn are consistent with Palma:2020ejf ; Fumagalli:2020nvq for the sharp turn case once the Hubble friction is ignored. For the broad turn case we revisit some of the assumptions made in Bjorkmo:2019qno and clarify the role of the Hubble friction terms in the derivation of the analytic solutions.
III.1 Sharp turns

### v3

@@ 2.1 Orthonormal frame and initial conditions
$\ddot{q}^{a}+\left(3H\delta^{a}_{~{}b}-2z^{a}_{~{}b}\right)\dot{q}^{b}+\left({k^{2}\over a^{2}}\delta^{a}_{~{}b}+M_{IJ}E^{I}_{~{}a}E^{J}_{~{}b}+z^{a}_{~{}c}z^{c}_{~{}b}-3Hz^{a}_{~{}b}-\dot{z}^{a}_{~{}b}\right)q^{b}=0\,.$ (2.16)
By an appropriate field redefinition, and using the Frenet-Serret equations, we can simplify the Lagrangian (2.15). By redefining the tangential component as $q_{\rm t}\equiv\sqrt{2\epsilon}\mathcal{R}$ and performing integrations by parts, the Lagrangian simplifies to
$L$ $=a^{3}\left\{{1\over 2}\dot{q}^{a}\dot{q}_{a}-{1\over 2a^{2}}\nabla q^{a}\nabla q_{a}-{\epsilon\over a^{2}}\nabla\mathcal{R}\nabla\mathcal{R}-z_{ab}\dot{q}^{a}q^{b}+{1\over 2}\omega^{2}q_{\rm n}^{2}+\epsilon\dot{\mathcal{R}}^{2}-2\omega\sqrt{2\epsilon}\dot{\mathcal{R}}q_{\rm n}\right.$ (2.17)
$\quad\left.-{1\over 2}\left(M_{ab}+z_{ac}z^{c}_{~{}b}\right)q^{a}q^{b}+\left(M_{\mathcal{R}{\rm b}}+\omega\tau_{\rm b}\right)\sqrt{2\epsilon}\mathcal{R}q_{\rm b}+\left(M_{\mathcal{R}{\rm n}}+3H\omega+\dot{\omega}+{\dot{\epsilon}H\over\epsilon}\right)\sqrt{2\epsilon}\mathcal{R}q_{\rm n}\right.$
$\quad\left.-\left[M_{\mathcal{R}\mathcal{R}}-\omega^{2}-{1\over 4}\eta^{2}H^{2}-{3\over 2}H\dot{\epsilon}-{1\over\epsilon}{\,\mathrm{d}\over\,\mathrm{d}t}\left({\epsilon\eta H\over 2}\right)\right]\epsilon\mathcal{R}^{2}\right\}\,,$
where now $a,b\neq\mathcal{R}$ and $\eta\equiv\dot{\epsilon}/(H\epsilon)$. Using the Frenet-Serret equations we find that the second and third lines which involve $M_{\mathcal{R}\mathcal{R}}$, $M_{\mathcal{R}\rm{n}}$ and $M_{\mathcal{R}\rm{b}}$ vanish identically. Thus, the quadratic Lagrangian for multi-field perturbations in terms of the curvature perturbation $\mathcal{R}$ and the isocurvature perturbations $q_{i}$ is
$L$ $=a^{3}\left[{1\over 2}\dot{q}^{i}\dot{q}_{i}-z_{ij}\dot{q}^{i}q^{j}-{1\over 2a^{2}}\nabla q^{i}\nabla q_{i}-{1\over 2}\left(M_{ij}+z_{im}z^{m}{}_{j}\right)q^{i}q^{j}+{1\over 2}\omega^{2}q_{\rm n}^{2}\right.$ (2.18)
$\quad\left.-2\omega\sqrt{2\epsilon}\dot{\mathcal{R}}q_{\rm n}-{\epsilon\over a^{2}}\nabla\mathcal{R}\nabla\mathcal{R}+\epsilon\dot{\mathcal{R}}^{2}\right]\,.$
Similar expressions for the second-order Lagrangian have been derived in [53, 54].
To properly assign initial conditions for the perturbations (e.g. Bunch-Davies) we need to transform the action into its canonical form. This can be done by using a transformation to another local orthonormal basis which can eliminate the kinetic couplings between the fields. We will work with the conformal time $\,\mathrm{d}\tau\equiv a^{-1}\,\mathrm{d}t$ and redefine the fields as $u^{i}\equiv aq^{i}$ to simplify the action (2.15) to
$S_{2}={1\over 2}\int\,\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\delta_{ij}{\,\mathrm{d}u^{i}\over\,\mathrm{d}\tau}{\,\mathrm{d}u^{j}\over\,\mathrm{d}\tau}-2aHZ_{ij}{\,\mathrm{d}u^{i}\over\,\mathrm{d}\tau}u^{j}-\delta_{ij}\nabla u^{i}\nabla u^{j}-(aH)^{2}m_{ij}u^{i}u^{j}\right]\,,$ (2.19)
@@
$\begin{split}Z_{ij}&\equiv\frac{z_{ij}}{H}\,,\\ \mathcal{M}_{ij}&\equiv\frac{M_{ij}}{H^{2}}\,,\\ m_{ij}&\equiv\mathcal{M}_{IJ}E_{i}^{~{}I}E_{j}^{~{}J}+Z_{il}Z^{l}_{~{}j}-\left(2-\epsilon\right)\delta_{ij}\,.\end{split}$ (2.20)
The expression (2.19) is not yet suitable for quantization due to the second term that mixes the fields and their derivatives. We can eliminate these terms by introducing explicit time dependence in the definition of the orthonormal basis vectors [49, 55, 56]. A time-dependent change of basis
$u^{i}=R^{i}_{~{}\alpha}(\tau)v^{\alpha}\,,$ (2.21)
@@
${\,\mathrm{d}R^{i}_{~{}\alpha}\over\,\mathrm{d}\tau}-aHZ^{i}_{~{}j}R^{j}_{~{}\alpha}=0\,,$ (2.22)
i.e. if the matrix is orthogonal. Regarding the initial condition of (2.22), we are completely free to choose any convenient, physically relevant time. In accordance with the previous literature we consider that the two basis vectors are aligned at horizon crossing $k\tau_{*}=-1$ and hence the initial condition becomes $R^{i}_{~{}\alpha}(\tau_{*})=\delta^{i}{}_{\alpha}$. In terms of the new fields we find a canonical second-order action as
$S_{2}={1\over 2}\int\,\mathrm{d}\tau\,\mathrm{d}^{3}x\left[\delta_{\alpha\beta}{\,\mathrm{d}v^{\alpha}\over\,\mathrm{d}\tau}{\,\mathrm{d}v^{\beta}\over\,\mathrm{d}\tau}-\delta_{\alpha\beta}\nabla v^{\alpha}\nabla v^{\beta}+(aH)^{2}(2-\epsilon)\delta_{\alpha\beta}v^{\alpha}v^{\beta}-(aH)^{2}\mathcal{M}_{ij}R^{i}_{~{}\alpha}R^{j}_{~{}\beta}v^{\alpha}v^{\beta}\right]\,,$ (2.23)
@@
$v^{\alpha}|_{\rm BD}={1\over\sqrt{2k}}e^{-ik\tau}\left(1-{i\over k\tau}\right)\,,\qquad\alpha=1,\cdots,\mathcal{N}\,,$ (2.25)
@@
$u^{i}|_{-k\tau\gg 1}=R^{i}_{~{}\alpha}(\tau,\tau_{*})v^{\alpha}|_{\rm BD}\,.$ (2.26)
2.2 Scale dependence of the power spectrum
At sufficiently late times we are interested in the scale dependence of the normalized power spectrum defined by
$\left\langle{\mathcal{R}}(\bm{k}){\mathcal{R}}(\bm{k}^{\prime})\right\rangle\equiv(2\pi)^{3}\delta^{(3)}\left(\bm{k}-\bm{k}^{\prime}\right){2\pi^{2}\over k^{3}}\mathcal{P}_{\mathcal{R}}\,.$ (2.27)
Due to the complexity of the coupled equations, in general one has to solve numerically the coupled equations of perturbations to find the superhorizon value of $\mathcal{R}$. However, we will show that when every background quantity, i.e. $\epsilon$, $\mathcal{M}_{\rm nn}\cdots$, is constant the scale dependence of the power spectrum can be read off the equations of perturbations without explicitly solving them. This can be done in the following way: First, in the differential equations for the independent variables of (2.24) we trade $\tau$ with [57]
$x\equiv-k\tau\,,$ (2.28)
@@ 2.2 Scale dependence of the power spectrum
$\tilde{v}\tilde{v}_{,x}^{*}-\tilde{v}^{*}\tilde{v}_{,x}=-i\,.$ (2.30)
@@
$\mathcal{R}_{\rm as}={H\over\sqrt{2\epsilon k^{3}}}xR_{~{}\alpha}^{1}\tilde{v}^{\alpha}_{\rm as}\,,$ (2.31)
@@
$(1-\epsilon)N=-\log x+\log\left(\frac{k}{a_{0}H_{0}}\right)\,,$ (2.33)
@@
$H=H_{0}x^{\epsilon/(1-\epsilon)}\left({k\over a_{0}H_{0}}\right)^{-\epsilon/(1-\epsilon)}\,,$ (2.34)
and the normalized power spectrum scales as $\mathcal{P}_{\mathcal{R}}\sim k^{-2\epsilon/(1-\epsilon)}$, irrespective of the number of fields. This remarkably has exactly the same form as the single-field power spectrum for scaling solutions (see e.g. [58] for the calculation in the single-field case).
@@
$\mathcal{P}_{\mathcal{R}}={H^{2}\over 4\pi^{2}}\frac{x^{2}|\tilde{u}_{\rm t}|^{2}}{\epsilon}\,.$ (2.35)
@@
$|\tilde{u}_{\rm t}|^{2}\approx\tilde{u}_{i}^{*}\tilde{u}_{i}=\tilde{v}_{\alpha}^{*}\tilde{v}_{\alpha}\,,$ (2.36)
and the final result does not depend on the rotation matrix $R$. If $\epsilon$ and $\eta$ vary slowly, but all other quantities are constant then the scale dependence of the power spectrum becomes identical to the single-field case to lowest order. [^3] This has been noticed before in cases where the orthogonal perturbation is heavy and can be integrated out, leading to an effective description of the curvature perturbation with a modified speed of sound [49]. However, our conclusion is not limited to heavy orthogonal fields. Otherwise, to find the asymptotic value of the curvature perturbation one has to solve the differential equations explicitly, which is the topic of section 3.2.
2.3 Hamiltonian description
@@
The canonical momenta for the Lagrangian (2.18) are
$\pi_{\mathcal{R}}$ $={\partial L\over\partial\dot{\mathcal{R}}}=2a^{3}\left(\epsilon\dot{\mathcal{R}}-\omega\sqrt{2\epsilon}q_{\rm n}\right)\,,$ (2.37)
$\pi_{i}$ $={\partial L\over\partial\dot{q}^{i}}=a^{3}\left(\dot{q}_{i}-z_{ij}q^{j}\right)\,,$ (2.38)
@@
$\mathcal{H}={1\over 2}{\pi^{i}\pi_{i}\over a^{3}}+z_{ij}\pi^{i}q^{j}+a{1\over 2}\nabla q^{i}\nabla q_{i}+{1\over 2}a^{3}M_{ij}q^{i}q^{j}+{3\over 2}a^{3}\omega^{2}q_{\rm n}^{2}+2{\omega\over\sqrt{2\epsilon}}\pi_{\mathcal{R}}q_{\rm n}+{1\over 4}{\pi_{\mathcal{R}}^{2}\over\epsilon a^{3}}+{\epsilon\over a^{2}}\nabla\mathcal{R}\nabla\mathcal{R}\,.$ (2.39)
@@
$\pi_{\mathcal{R}}$ $=a^{3}\sqrt{2\epsilon}\left(\,\text{D}_{t}q_{\rm t}-\frac{1}{2}H\eta q_{\rm t}-\omega q_{\rm n}\right)\,,$ (2.40)
$\pi_{i}$ $=a^{3}\left(\,\text{D}_{t}q_{i}-\omega q_{\rm n}\delta_{i{\rm n}}\right)\,.$ (2.41)
@@
$\dot{q}^{a}={\partial\mathcal{H}\over\partial\pi_{a}}\quad\text{and}\quad\dot{\pi}_{a}=-{\partial\mathcal{H}\over\partial q^{a}}$ (2.42)
@@
$\dot{\mathcal{R}}$ $={1\over 2}{\pi_{\mathcal{R}}\over\epsilon a^{3}}+\sqrt{\frac{2}{\epsilon}}\omega q_{\rm n}\,,$ (2.43)
$\dot{\pi}_{\mathcal{R}}$ $=2\epsilon a\nabla^{2}\mathcal{R}\,,$ (2.44)
$\dot{q}_{i}$ $={\pi_{i}\over a^{3}}+z_{ij}q^{j}\,,$ (2.45)
$\dot{\pi}_{i}$ $=-2{\omega\over\sqrt{2\epsilon}}\pi_{\mathcal{R}}\delta_{i{\rm n}}+z_{ij}\pi^{j}+a^{3}\left({1\over a^{2}}\nabla^{2}q_{i}-M_{ij}q^{j}-3\omega^{2}\delta_{i{\rm n}}q_{\rm n}\right)\,.$ (2.46)
@@
$P_{\cal R}\equiv\frac{\pi_{\cal R}}{2\epsilon a^{3}H}\quad\text{and}\quad P_{i}\equiv\frac{\pi_{i}}{a^{3}H}\,,$ (2.47)
@@
$\mathcal{R}^{\prime}$ $=P_{\mathcal{R}}+\sqrt{\frac{2}{\epsilon}}\Omega q_{\rm n}\,,$ (2.48)
$P_{\mathcal{R}}^{\prime}$ $=-(3-\epsilon+\eta)P_{\mathcal{R}}-\kappa^{2}\mathcal{R}\,,$ (2.49)
$q_{i}^{\prime}$ $=P_{i}+Z_{ij}q^{j}\,,$ (2.50)
$P_{i}^{\prime}$ $=-(3-\epsilon)P_{i}-2\sqrt{2\epsilon}\Omega P_{\mathcal{R}}\delta_{i{\rm n}}+Z_{ij}P^{j}-\kappa^{2}q_{i}-\mathcal{M}_{ij}q^{j}-3\Omega^{2}\delta_{i{\rm n}}q_{\rm n}\,,$ (2.51)
@@
$\Omega\equiv{\omega\over H}\quad\text{and}\quad\kappa\equiv{k\over aH}\,.$ (2.52)
The formulation in terms of the fields and their “momenta” [^4] Even though we will keep referring to the variables $P_{i}$ as momenta, these are not the canonical momenta of the problem. nicely illustrates the superhorizon sourcing of the curvature perturbation due to a non-vanishing normal perturbation [50]: Excluding runaway solutions $\mathcal{R}\propto a^{2}$ which would render the perturbation theory invalid, (2.49) implies that for $\kappa\rightarrow 0$ the momentum $P_{\mathcal{R}}$ is exponentially decaying. Therefore, $\mathcal{R}$ is sourced by the normal perturbation outside the horizon, whereas $q_{\rm n}$ is not. Moreover, for a sudden change of the bending parameters the previous set of equations has no singularities and no special junction conditions are needed at the discontinuities (as was done in e.g. [38, 39, 47, 45]) apart from continuity of the fields and their momenta.
3 Analytic solutions for constant turns
To gain further physical insights, in this section we consider a simple two-field system that undergoes a sharp or broad turn. We can find a set of analytic solutions of perturbations, and in particular can identify how the positive eigenvalues enhance the perturbations. Note that while we use Hamilton’s equations to study the behaviour of the two perturbations, we have checked that our results for a single turn are consistent with [38, 39] for the sharp turn case once the Hubble friction is ignored. For the broad turn case we revisit some of the assumptions made in [59] and clarify the role of the Hubble friction terms in the derivation of the analytic solutions.
3.1 Sharp turns